5x^2 - y, when x = 4 and y = 24
I got 56 as my answer.

Answer by ikleyn(52833)

.

5x^2 - y,  when x = 4  and  y = 24,  is


    5*(4^2) - 24 = 5*16 - 24 = 80 - 24 = 56.    ANSWER


Your answer is correct.

Solved.